OpenStudy (anonymous):

Find the indicated derivative and simplify d/dx[7x^3-3x^2/3squarerootx^2]

OpenStudy (mimi_x3):

Is it: \[\large \frac{7x^{3}-3x^{2}}{3\sqrt{x^{2}}} \]?

OpenStudy (campbell_st):

simplfy the problem by using the index laws \[7x^3/3(^2)^{-12/} - 3x^2/3(x^2)^{-1/2}\] simplifies to \[7/3x^2 - x\] its now a bit easier to find dy/dx

OpenStudy (campbell_st):

oops got it wrong it should be simplified to \[7x^2/3 - x\]

OpenStudy (anonymous):

use the formula dy/dx=u/v,,,,,,,=v(d/dx(u))-u(d/dx(v))all over v^2 where y respect to x